# Model-free High Dimensional Mediator Selection with False Discovery Rate Control

### v1

@@ 1 Introduction
With our ADNI structural MRI data as candidate mediators, the existing methods present limitations. First, path-a effects are typically small due to subtle gender differences in brain structure and function compared to the large within-gender variability. Second, interactions between various structural MRI measures potentially influence cognitive outcomes. For example, a study found that in individuals with lower gray matter (GM) volumes, increased functional connectivity between the medial prefrontal cortex and the lateral intraparietal cortex was associated with better memory performance. However, this association was not found in individuals with higher GM volumes. This means that GM volume might change how functional connectivity and memory are related (He et al., 2012). Furthermore, another study found that the baseline white matter hyperintensity volume (WMHV) worked with age to affect CDR-SB scores in females but not in males, showing that brain structure changes might affect memory loss differently in men and women (Schweitzer et al., 2024). Therefore, a more advanced nonlinear mediation framework is required to capture the true mechanistic pathways in AD progression.
@@ 2.1 Model settings
[…]
@@
[…]
@@
[…]
@@
is for two-way interaction terms and $\bm{\delta}$ is its coefficient. Our method also allows for general nonlinear model forms of path-b to accommodate various types of response variables, including categorical and count data. For example, we can have generalized additive models like $g^{-1}(\mathbb{E}[Y])=f_{1}(X)+\sum_{j=1}^{p}f_{2j}(M_{j})+f_{3}(\mathbf{V})+\epsilon_{2}$ or a nonparametric model $Y=f(X,\mathbf{M},\mathbf{Z},\mathbf{V},\epsilon_{2})$ where $f_{1}$, $f_{2j}$, $f_{3}$, $f$ are unknown functions.
Under our general model setting, we frame the mediator selection problem as a multiple testing problem, where for each $j\in\{1,\cdots,p\}$, $H_{0j}=H_{aj}\cup H_{bj}$ is a union null hypothesis, and =$\mathcal{H}=\{k\in[p]:H_{0k}~{}\text{is true}\}$. Our goal is to identify a selection set $\widehat{S}\subset[p]$ and control for the FDR defined as
@@ 2.2 Preliminaries
$f([\mathbf{Z}^{1},\widetilde{\mathbf{Z}}^{1}],\ldots,[\mathbf{Z}^{k},\widetilde{\mathbf{Z}}^{k}]_{\text{Swap}(S)},\ldots,[\mathbf{Z}^{K},\widetilde{\mathbf{Z}}^{K}])=f([\mathbf{Z}^{1},\widetilde{\mathbf{Z}}^{1}],\ldots,[\mathbf{Z}^{k},\widetilde{\mathbf{Z}}^{k}],\ldots,[\mathbf{Z}^{K},\widetilde{\mathbf{Z}}^{K}])\odot\epsilon(S),$
@@ 1.1 Path-a knockoff generation
    For path-a, for each $j\in[p]$, we first generate knockoff $\widetilde{M}_{j}$ for $M_{j}$, where $\widetilde{M}_{j}$ is independent of $X$ and has identical distribution with $M_{j}$ conditioning on $\mathbf{V}$. We can generate knockoff $(\widetilde{M}_{j},\widetilde{\mathbf{V}})$ for $(M_{j},\mathbf{V})$ using the Model-X knockoff construction approach (Candès et al., 2018). Depending on the specific data type of $M_{j}$, we can use the Gaussian approach, second-order approach, sequential approach, or even deep learning-based approaches (Kormaksson et al., 2021; Romano et al., 2020) for the knockoff construction.
@@ 1.2 Path-a test statistics computation
  • Step 2: Construction of knockoff based test-statistics for path-b. For path-b, we will construct knockoff $\widetilde{\mathbf{M}}^{b}$ for $\mathbf{M}$ conditioning on $\mathbf{V}$ and $X$. This can be done through generating fixed or Model-X knockoff of $(\mathbf{M},\mathbf{V},X)$ as $(\widetilde{\mathbf{M}}^{b},\widetilde{\mathbf{V}}^{b},\widetilde{X}^{b})$. Denote $\widetilde{\mathbf{M}}^{b}=(\widetilde{M}_{1},\widetilde{M}_{2},\cdots,\widetilde{M}_{p})$. We then fit penalized regression models or non-parametric regression models (e.g. random forests or a deep learning model) for $Y$ on $(\mathbf{M},\widetilde{\mathbf{M}}^{b},\mathbf{V},X)$ and use the absolute value of coefficients or certain importance measure of variables (e.g. importance score or SHAP value) for the variables $M_{j}$, $\widetilde{M}_{j}$ as $Z^{b}_{j}$ and $\widetilde{Z}^{b}_{j}$. Denote $(\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b})=(Z^{b}_{1},\cdots,Z^{b}_{p},\widetilde{Z}^{b}_{1},\cdots,\widetilde{Z}^{b}_{p})$, which can be shown to satisfy the requirement that for any $S\in\mathcal{H}_{b}$, $(\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b})\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}(\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b})_{Swap(S)}$ (Candès et al., 2018).
@@
[…]
@@
[…]
@@ 2.4 Theoretical results
The proof of Theorem 1 is postponed to Web Appendix A.1.
@@ 3.1 Data Generation and Simulation Settings
For each subject $i\in[n]$, the exposure variable $X_{i}\in\mathbb{R}$ and the covariate $V_{i}\in\mathbb{R}$ are generated from a multivariate normal distribution, with mean $\begin{pmatrix}0\\ 0\end{pmatrix}$ and variance-covariance matrix $\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}$. The correlated potential mediators $\mathbf{M}_{i}=(M_{i1},\dots,M_{ip})^{\top}$ is generated using $\mathbf{M}_{i}=X_{i}\bm{\alpha}+V_{i}\bm{\eta}_{1}+\bm{\epsilon}_{1i}$, where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{p})^{\top}$,$\bm{\eta_{1}}=(\eta_{11},\dots,\eta_{1p})^{\top}$ and $\bm{\epsilon}_{1i}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma})$, where $\bm{\Sigma}$ is a variance-covariance matrix with compound symmetry structure (i.e.,diagonal entries are 1 and off-diagonal entries are $\rho$). The outcome $Y_{i}$ is generated as $Y_{i}=X_{i}\gamma+\mathbf{M}^{\top}\bm{\beta}+V_{i}\eta_{2}+\epsilon_{2i}$, where $\bm{\epsilon}_{2i}\sim\mathcal{N}(0,\sigma^{2})$. Detailed simulation settings for parameters $\bm{\alpha},\bm{\eta}_{1},\rho,\gamma,\bm{\beta},\eta_{2},\sigma$ are provided in Appendix C.1.
@@
Further details on the data generation process are provided in Appendix C.1.
@@ 3.2 Method specification
For our proposed method, for path-a, we use the Guassian Model-X method to generate knockoffs $(\widetilde{M}_{j},\widetilde{V})$ for $(M_{j},V)$, and fit a regression model of the form $X\sim(M_{j},\widetilde{M}_{j},V)$ to obtain test statistics $(\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a})$ using the absolute value of the coefficients (more details are in Apendix A). For path-b linear settings, we use the PLS option in Remark 1 to derive the test statistics $(\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b})$. For the nonlinear setting, we fit random forests models and extract variable importance scores to construct $(\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b})$. For HIMA methods, (i.e.,HIMA, HIMA2 and eHIMA), we use the HIMA package version 2.2.2. with their default methods. We set the controlled FDR level $q=0.2$ for our proposed method and all HIMA methods. For method MediationFDR, we apply the MediationFDR package version 1.0.0 with the default methods and to control the overall FDR at $0.2$, we set $q_{1}=0.1$ in step 1 and $q_{2}=0.1$ in step 4. For the method MCP, we use the package MultiMed version 2.24.0 with the default methods and set the thresholds as 0.1 for both path-a and path-b.
@@ 4 Data Analysis
To pre-process the structural MRI data, we first convert ADNI dataset to follow the Brain Imaging Data Structure (BIDS) rule via Clinica (Routier et al., 2021) and applied fmriPrep (Esteban et al., 2019) to preprocess the data. This pipeline outputs 555 structural MRI-derived measures of each subjects are obtained after we apply fmriPrep, such as cortical thickness, surface area, and brain volumes, each corresponding to specific brain regions or morphometric features, such as the average cortical thickness in various gyri and sulci, and volumetric data for subcortical structures. The full list of 555 structural MRI-derived measures can be found in Web Appendix D.
The outcome of interest in this study is the change in multiple cognitive assessment scores from baseline to 17 years of follow-up. These assessments are typically conducted at baseline and repeated at regular intervals (e.g., every 3 or 6 months). To account for the missing visits, we apply the last observation carried forward (LOCF) method (Lachin, 2016) to impute missing cognitive scores. To evaluate different aspects of cognitive function, we apply multiple cognitive scores, including the Clinical Dementia Rating Sum of Boxes (CDR-SB), Alzheimer’s Disease Assessment Scale-Cognitive Subscale (ADAS-Cog) and Mini-Mental State Examination (MMSE). Higher CDR-SB and ADAS-Cog scores, and lower MMSE scores, are each indicative of greater Alzheimer’s disease severity. More details about those cognitive scores can be found in Appendix D.
Finally, we apply our proposed GKMS method to identify important structural MRI-derived measures from a set of $p=555$ candidate correlated mediators that may mediate the relationship between gender and changes in cognitive scores among $n=1097$ MCI patients. We consider Age at baseline as a confounding variable $V$. The FDR is set as $q=0.2$. For exposure X, we set males as the reference group. $\mathbf{Y}$ are the changes from baselines in the cognitive assessment scores CDR-SB, ADAS-Cog and MMSE. Both the outcome $\mathbf{Y}$ and the mediators $\mathbf{M}$ are scaled without centering ensuring that the values the $\alpha$ and $\beta$ are the standardized regression coefficients. Tables 3, 4 and 5 summarized all the mediator selection results with their estimated mediation effect for changes in CDR-SB, ADAS-Cog and MMSE scores, respectively. For our proposed method with random forest estimation in path-b (GKMS (RF)), we evaluate the mediator effect using g-formula with numerical Monte Carlo integration as described by Imai et al. (2010), For the other methods, we employed structural equation modeling (SEM) through the lavaan R package (Rosseel, 2012). More details on the data analyses are provided in Appendix D.
@@
Method Selected Mediator Effect (95% CI; p-value)
GKMS (RF)* Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; $<$0.001)
$\beta$ = -0.157 (-0.388, -0.228; $<$0.001))
Mediation = 0.063 (0.013, 0.092; $<$0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = -0.213 (-0.458, -0.293;$<$0.001)
Mediation =0.092 (0.051, 0.137; $<$0.001)
Volume of Hippocampus (R) $\alpha$ = -0.551 (-0.658, -0.444; $<$0.001)
$\beta$ = -0.162 (-0.393, -0.236; $<$0.001)
Mediation = 0.026 (0.004, 0.072; 0.02)
Thick Avg of Angular Gyrus (L) $\alpha$ = 0.23 (0.115, 0.346; $<$0.001)
$\beta$ = -0.135 (-0.342, -0.201; $<$0.001)
Mediation = -0.024 (-0.045, -0.005; $<$0.001)
ThickAvg of Sulcus (L) $\alpha$ = 0.192 (0.082, 0.302; $<$0.001)
$\beta$ = -0.165 (-0.405, -0.234; $<$0.001)
Mediation =-0.016 (-0.041, 0.003; 0.08)
GKMS (Lasso) Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; $<$0.001)
$\beta$ = -0.132 (-0.193; -0.071;$<$0.001)
Mediation = 0.069 (0.034, 0.104; $<$0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = -0.178 (-0.238; -0.118;$<$0.001)
Mediation = 0.09 (0.054, 0.126; $<$0.001)
Volume of Amygdala (R) $\alpha$ = -0.596 (-0.706, -0.487; $<$0.001)
$\beta$ = -0.111 (-0.171; -0.052;$<$0.001)
Mediation = 0.066 (0.029, 0.104; 0.001)
Volume of Inf.Lat.Vent1 (R) $\alpha$ = -0.483(-0.591, -0.376; $<$0.001)
$\beta$ = 0.15 (0.089; 0.211;$<$0.001)
Mediation = -0.072 (-0.106, -0.039; $<$0.001)
Volume of Brain Stem $\alpha$ = -0.955(-1.061, -0.848; $<$0.001)
$\beta$ = -0.121 (0.06; 0.183;$<$0.001)
Mediation = -0.116 (-0.176, -0.056; $<$0.001)
MediationFDR Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = -0.376 (-0.438; -0.314;$<$0.001)
Mediation = 0.19 (0.139, 0.241; $<$0.001)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = -0.282 (-0.343, -0.222; $<$0.001)
Mediation = 0.143 (0.099, 0.186; $<$0.001)
Volume of Inf.Lat.Vent (R) $\alpha$ = -0.483(-0.591, -0.376; $<$0.001)
$\beta$ = 0.135 (0.074, 0.196; $<$0.001)
Mediation = -0.065 (-0.098, -0.032; $<$0.001)
ThickAvg of Parahip2 (R) $\alpha$ = 0.123 (0.007, 0.24; 0.037)
$\beta$ = -0.128 (-0.185, -0.072; $<$0.001)
Mediation = -0.016 (-0.032, 0.001; 0.059)
eHIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = -0.376 (-0.438; -0.314;$<$0.001)
Mediation = 0.19 (0.139, 0.241; $<$0.001)
  • 1 Inf.Lat.Vent refers to the Inferior Lateral Ventricle.
Method Selected Mediator Effect (95% CI; p-value)
GKMS (RF)* Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; $<$0.001)
$\beta$ = -0.118 (-0.364, -0.186; $<$0.001)
Mediation = 0.041 (0.01, 0.073; 0.016)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = -0.171 (-0.411, -0.249; $<$0.001)
Mediation = 0.102 (0.039, 0.127; $<$0.001)
ThickAvg of Angular Gyrus(L) $\alpha$ = 0.233 (0.118, 0.349; $<$0.001)
$\beta$ = -0.164 (-0.418, -0.222 $<$0.001)
Mediation = -0.026 (-0.057, -0.005; 0.02)
$\beta$ = -0.177 (-0.422, -0.235; $<$0.001)
Mediation =-0.004 (-0.032, 0.008; 0.22)
Volume of white matter in ITL 2(L) $\alpha$ = -0.895 (-1.004, -0.785; $<$0.001)
$\beta$ = 0.006 (-0.155, 0.185; 0.9)
Mediation =-0.026 (-0.057, 0.022; 0.34)
GKMS (Lasso) Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; $<$0.001)
$\beta$ = -0.083 (-0.145, -0.02; 0.009)
Mediation = 0.043 (0.009, 0.077; 0.012)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = -0.139 (-0.201, -0.078; $<$0.001)
Mediation = 0.07 (0.036, 0.105; $<$0.001)
ThickAvg of Angular Gyrus(L) $\alpha$ = 0.233 (0.118, 0.349; $<$0.001)
$\beta$ = -0.099 (-0.157, -0.041; 0.001)
Mediation = -0.023 (-0.041, -0.005; 0.011)
ThickAvg of S.temporal.sup(R) $\alpha$ = 0.092 (-0.023, 0.207; 0.117)
$\beta$ = -0.167 (-0.226, -0.109; $<$0.001)
Mediation = -0.015 (-0.035, 0.005; 0.131)
Volume of Choroid Plexus (L) $\alpha$ = -0.719 (-0.826, -0.611; $<$0.001)
$\beta$ = 0.081 (0.019, 0.143; 0.011)
Mediation = -0.058 (-0.104, -0.013; 0.012)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = -0.324 (-0.387, -0.261; $<$0.001)
Mediation = 0.164 (0.116, 0.211; $<$0.001)
eHIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = -0.324 (-0.387, -0.261; $<$0.001)
Mediation = 0.164 (0.116, 0.211; $<$0.001)
$\beta$ = -0.063 (-0.123, -0.004; 0.036)
Mediation = 0.013 (-0.001, 0.027; 0.078)
ThickAvg of Pole temporal(R) $\alpha$ = 0.259 (0.143, 0.376; $<$0.001)
$\beta$ = -0.084 (-0.146, -0.023; 0.007)
Mediation = -0.022 (-0.041, -0.003; 0.022)
  • We use version ADAS-13 representing the results of ADAS-Cog.
Method Selected Mediator Effect (95% CI; p-value)
GKMS (RF)* Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; $<$0.001)
$\beta$ = 0.129 (0.184, 0.375; $<$0.001)
Mediation = -0.063 (-0.109, -0.016; 0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ =0.165 (0.256, 0.409; $<$0.001)
Mediation = -0.077 (-0.142, -0.026; $<$0.001)
Volume of Amygdala (R) $\alpha$ = -0.596 (-0.706, -0.487; $<$0.001)
$\beta$ = 0.157 (0.224, 0.396; $<$0.001)
Mediation =-0.036 (-0.075, 0.001; 0.04)
Volume of Hippocampus (R) $\alpha$ = -0.551 (-0.658, -0.444; $<$0.001)
$\beta$ = 0.132 (0.2, 0.366; $<$0.001)
Mediation = -0.076 (-0.122, -0.029; 0.032)
ThickAvg of G-front-middle1 (L) $\alpha$ = 0.29 (0.175, 0.404; $<$0.001)
$\beta$ = 0.136 (0.206, 0.392; $<$0.001)
Mediation = 0.028 (0.013, 0.056; $<$0.001)
GKMS (Lasso) Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; $<$0.001)
$\beta$ = 0.111 (0.048, 0.174; 0.001)
Mediation = -0.058 (-0.093, -0.023; 0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = 0.184 (0.122, 0.246; $<$0.001)
Mediation = -0.093 (-0.13, -0.056; $<$0.001)
Volume of Amygdala (R) $\alpha$ = -0.596 (-0.706, -0.487; $<$0.001)
$\beta$ = 0.068 (0.007, 0.129; 0.028)
Mediation = -0.041 (-0.078, -0.004; 0.032)
Volume of Inf.Lat.Vent2 (R) $\alpha$ = -0.483(-0.591, -0.376; $<$0.001)
$\beta$ = -0.088 (-0.15, -0.026; 0.006)
Mediation = 0.042 (0.011, 0.074; 0.008)
Volume of Medial Orbitofrontal (L) $\alpha$ = -0.702 (-0.816, -0.588; $<$0.001)
$\beta$ = -0.147 (-0.206, -0.088; $<$0.001)
Mediation = 0.103 (0.059, 0.148; $<$0.001)
MediationFDR Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = 0.336 (0.273, 0.399; $<$0.001)
Mediation = -0.17 (-0.218, -0.121; $<$0.001)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = 0.219 (0.158, 0.281; $<$0.001)
Mediation = -0.111 (-0.15, -0.072; $<$0.001)
ThickAvg of S.temporal.sup3 (L) $\alpha$ = 0.12 (0.007, 0.233; 0.037)
$\beta$ = 0.246 (0.187, 0.305; 0)
Mediation = 0.03 (0.001, 0.058; 0.043)
eHIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; $<$0.001)
$\beta$ = 0.336 (0.273, 0.399; $<$0.001)
Mediation = -0.17 (-0.218, -0.121; $<$0.001)
  • 2 Inf.Lat.Vent refers to the Inferior Lateral Ventricle.

### v2

@@ 1 Introduction
With the ADNI structural MRI data as candidate mediators, the existing methods present limitations. First, path-a effects are typically small due to subtle gender differences in brain structure and function compared to the large within-gender variability. Second, interactions between various structural MRI measures potentially influence cognitive outcomes. For example, a study found that in individuals with lower gray matter (GM) volumes, increased functional connectivity between the medial prefrontal cortex and the lateral intraparietal cortex was associated with better memory performance. However, this association was not found in individuals with higher GM volumes. This means that GM volume might change how functional connectivity and memory are related (He et al., 2012). Furthermore, another study found that the baseline white matter hyperintensity volume (WMHV) worked with age to affect CDR-SB scores in females but not in males, showing that brain structure changes might affect memory loss differently in men and women (Schweitzer et al., 2024). Therefore, a more advanced nonlinear mediation framework is required to capture the true mechanistic pathways in AD progression.
@@ 2.1 Model settings
$\text{path-a:}~~H_{aj}:M_{j}\perp\!\!\!\perp X|\mathbf{V},$
$\text{path-b:}~~H_{bj}:Y\perp\!\!\!\perp M_{j}|M_{-j},X,\mathbf{V},$
$\text{where $M_{-j}:=\{M_{k}:k\in[p],k\neq j\}$}.$ (1)
@@
$\text{path-a:}\quad M_{j}=c_{1j}+X\alpha_{j}+\mathbf{V}^{\top}\bm{\eta}_{1j}+\epsilon_{1j},\quad\textnormal{for}~j\in[p],$ (2)
@@
$\text{path-b:}\quad Y=c_{2}+X\gamma+\mathbf{M}^{\top}\bm{\beta}+\mathbf{V}^{\top}\bm{\eta}_{2}+\epsilon_{2},$ (3)
@@
is for two-way interaction terms and $\bm{\delta}$ is its coefficient.
Our method also allows for general nonlinear model forms of path-b to accommodate various types of response variables, including categorical and count data. For example, we can have generalized additive models like $g^{-1}(\mathbb{E}[Y])=f_{1}(X)+\sum_{j=1}^{p}f_{2j}(M_{j})+f_{3}(\mathbf{V})+\epsilon_{2}$ or a nonparametric model $Y=f(X,\mathbf{M},\mathbf{Z},\mathbf{V},\epsilon_{2})$ where $f_{1}$, $f_{2j}$, $_{3}, ff$ are unknown functions.
Under our general model setting, we frame the mediator selection problem as a multiple testing problem, where for each $j\in\{1,\cdots,p\}$, $H_{0j}=H_{aj}\cup H_{bj}$ is a union null hypothesis, and $\mathcal{H}=\{k\in[p]:H_{0k}~\text{is true}\}$. Our goal is to identify a selection set $\widehat{S}\subset[p]$ and control for the FDR defined as
@@ 2.2 Preliminaries
$f($ $[\mathbf{Z}^{1},\widetilde{\mathbf{Z}}^{1}],\ldots,[\mathbf{Z}^{k},\widetilde{\mathbf{Z}}^{k}]_{\text{Swap}(S)},\ldots,[\mathbf{Z}^{K},\widetilde{\mathbf{Z}}^{K}])$
$=f([\mathbf{Z}^{1},\widetilde{\mathbf{Z}}^{1}],\ldots,[\mathbf{Z}^{k},\widetilde{\mathbf{Z}}^{k}],\ldots,[\mathbf{Z}^{K},\widetilde{\mathbf{Z}}^{K}])\odot\epsilon(S),$
@@ 1.1 Path-a knockoff generation
[…]
@@ 1.2 Path-a test statistics computation
  • Step 2: Construction of knockoff based test-statistics for path-b. For path-b, we will construct knockoff $\widetilde{\mathbf{M}}^{b}$ for $\mathbf{M}$ conditioning on $\mathbf{V}$ and $X$. This can be done through generating fixed or Model-X knockoff of $(\mathbf{M},\mathbf{V},X)$ as $(\widetilde{\mathbf{M}}^{b},\widetilde{\mathbf{V}}^{b},\widetilde{X}^{b})$. Denote $\widetilde{\mathbf{M}}^{b}=(\widetilde{M}_{1},\widetilde{M}_{2},\cdots,\widetilde{M}_{p})$. We then fit penalized regression models or non-parametric regression models (e.g. random forests or a deep learning model) for $Y$ on $(\mathbf{M},\widetilde{\mathbf{M}}^{b},\mathbf{V},X)$ and use the absolute value of coefficients or certain importance measure of variables (e.g. importance score or SHAP value) for the variables $M_{j}$, $\widetilde{M}_{j}$ as $Z^{b}_{j}$ and $\widetilde{Z}^{b}_{j}$. Denote $(\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b})=(Z^{b}_{1},\cdots,Z^{b}_{p},\widetilde{Z}^{b}_{1},\cdots,\widetilde{Z}^{b}_{p})$, which can be shown to satisfy the requirement that for any $S\in\mathcal{H}_{b}$, $(\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b})\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}(\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b})_{Swap(S)}$ (Candès et al., 2018). Further details on the test statistics compatible with each knockoff construction method are provided in Web Appendix A.2.
@@
    $\tau=\min\left\{t>\mathcal{W}_{+}:\frac{\#\{j:W_{j}\leq-t\}}{\#\{j:W_{j}\geq t\}\vee 1}\leq q\right\}~~\textnormal{(Knockoff)},$
@@
    $\tau+=\min\left\{t>\mathcal{W}_{+}:\frac{1+\#\{j:W_{j}\leq-t\}}{\#\{j:W_{j}\geq t\}\vee 1}\leq q\right\}~~\textnormal{(Knockoff+)},$
@@
2.4 Quantification of the mediation effects
After selecting the set of mediators, we further make inference on the natural direct effect (NDE) and natural indirect effect (NIE) under our general model settings.
[…]
For the general nonparametric model cases, under the counterfactual framework, denote $M_{k}(x)$ for $k\in[p]$ as the $k$th potential mediator if the exposure value is set as $x$; and $\mathbf{M}(x)\in\mathbb{R}^{p}$ as the potential mediators if the exposure value is set as $x$, i.e. $\mathbf{M}(x)=(M_{1}(x),\cdots,M_{p}(x))$. For the outcome $Y$, define $Y(x,M_{1}(x_{1}^{\prime}),\cdots,M_{p}(x_{p}^{\prime}))$ as the potential outcome value when the exposure is set at $x$ and mediators are set as $(M_{1}(x_{1}^{\prime}),\cdots,M_{p}(x_{p}^{\prime}))$.
We have
NDE $=\mathbb{E}\left[Y(1,\mathbf{M}(0))-Y(0,\mathbf{M}(0))\right],$ (7)
NIE $=\mathbb{E}\left[Y(1,\mathbf{M}(1))-Y(1,\mathbf{M}(0))\right].$ (8)
Following Imai et al. (2010), we implement a Monte Carlo integration algorithm to evaluate the integration of $\widehat{\mathbb{E}}\left[Y\left(x,M_{1}(x_{1}),\dots,M_{p}(x_{p})\right)\right]$ numerically.
  • Step 1: Sample $(M_{1}(x_{1}),\ldots,M_{p}(x_{p}))$ for different values of $(x_{1},\ldots,x_{p})\in\{0,1\}^{p}$ for all $\bm{V}_{i}$;
  • Step 2: Sample $Y(x,M_{1}(x_{1}),\ldots,M_{p}(x_{p}))$ for different values of $x\in\{0,1\}$ and $(x_{1},\ldots,x_{p})\in\{0,1\}^{p}$ for all $\bm{V}_{i}$;
  • Step 3: Compute $\widehat{\mathbb{E}}\left[Y(x,M_{1}(x_{1}),\ldots,M_{p}(x_{p}))\right]$ by averaging sampled $Y(x,M_{1}(x_{1}),\ldots,M_{p}(x_{p}))$ and average over individual $i$. Then compute $\widehat{\text{NDE}},\widehat{\text{NIE}}$, and $\widehat{\text{NIE}}_{k}$, where
    $\widehat{\text{NDE}}=\widehat{\mathbb{E}}\left[Y(1,\mathbf{M}(0))-Y(0,\mathbf{M}(0))\right],$
    $\widehat{\text{NIE}}=\widehat{\mathbb{E}}\left[Y(1,\mathbf{M}(1))-Y(1,\mathbf{M}(0))\right],\text{ and}$
    $\widehat{\text{NIE}}_{k}=\widehat{\mathbb{E}}\left[Y\left(0,\mathbf{M}_{1:k-1}(0),M_{k}(1),\mathbf{M}_{k+1:p}(0)\right)\right]-\widehat{\mathbb{E}}\left[Y\left(0,\mathbf{M}(0)\right)\right],$
    where $\mathbf{M}_{1:k-1}(0)=(M_{1}(0),\cdots,M_{k-1}(0))$, and $\mathbf{M}_{k+1:p}(0)=(M_{k+1}(0),\cdots,M_{p}(0))$.
@@ 2.5 Theoretical results
The proof of Theorem 1 is postponed to Web Appendix B.
@@ 3.1 Data Generation and Simulation Settings
For each subject $i\in[n]$, the exposure variable $X_{i}\in\mathbb{R}$ and the covariate $V_{i}\in\mathbb{R}$ are generated from a multivariate normal distribution, with mean $\begin{pmatrix}0\\ 0\end{pmatrix}$ and variance-covariance matrix $\begin{pmatrix}1&0.5\\ 0.5&1\end{pmatrix}$. The correlated potential mediators $\mathbf{M}_{i}=(M_{i1},\dots,M_{ip})^{\top}$ is generated using $\mathbf{M}_{i}=X_{i}\bm{\alpha}+V_{i}\bm{\eta}_{1}+\bm{\epsilon}_{1i}$, where $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{p})^{\top}$$\bm{\eta_{1}}=(\eta_{11},\dots,\eta_{1p})^{\top}$ and $\bm{\epsilon}_{1i}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma})$, where $\bm{\Sigma}$ is a variance-covariance matrix with compound symmetry structure (i.e.,diagonal entries are 1 and off-diagonal entries are $\rho$). The outcome $Y_{i}$ is generated as $Y_{i}=X_{i}\gamma+\mathbf{M}^{\top}\bm{\beta}+V_{i}\eta_{2}+\epsilon_{2i}$, where $\bm{\epsilon}_{2i}\sim\mathcal{N}(0,\sigma^{2})$. Detailed simulation settings for parameters $\bm{\alpha},\bm{\eta}_{1},\rho,\gamma,\bm{\beta},\eta_{2},\sigma$ are provided in Appendix C.1.
@@
Further details on the data generation process are also provided in Appendix C.1.
@@ 3.2 Method specification
[…]
@@ 4 Data Analysis
To obtain the structural MRI data, we first convert ADNI dataset to follow the Brain Imaging Data Structure (BIDS) rule via Clinica (Routier et al., 2021) and applied fMRIPrep (Esteban et al., 2019) with the FreeSurfer (fischl2012freesurfer) reconstruction workflow to pre-process the data. This pipeline outputs 555 structural MRI-derived measures of each subjects (e.g. cortical thickness, surface area and brain volumes) corresponding to specific brain regions or morphometric features, such as the average cortical thickness in various gyri and sulci, and volumetric data for subcortical structures. The full list of the 555 structural MRI-derived measures can be found in Web Appendix D.1.
The outcome of interest in this study is the change in multiple cognitive assessment scores from baseline to 17 years of follow-up. These assessments are typically conducted at baseline and repeated at regular intervals (e.g., every 3 or 6 months). To account for the missing visits, we apply the last observation carried forward (LOCF) method (Lachin, 2016) to impute missing cognitive scores. To evaluate different aspects of cognitive function, we apply multiple cognitive scores, including the Clinical Dementia Rating Sum of Boxes (CDR-SB), Alzheimer’s Disease Assessment Scale-Cognitive Subscale (ADAS-Cog) and Mini-Mental State Examination (MMSE). Higher CDR-SB and ADAS-Cog scores, and lower MMSE scores, are each indicative of greater Alzheimer’s disease severity. More details about those cognitive scores can be found in Appendix D.2.
Finally, we apply our proposed GKMS method to identify important structural MRI-derived measures from a set of $p=555$ candidate correlated mediators that may mediate the relationship between gender and changes in cognitive scores among $n=1097$ MCI patients. We consider Age at baseline as a confounding variable $V$. The FDR is set as $q=0.2$. For exposure X, we set males as the reference group. $\mathbf{Y}$ are the changes from baselines in the cognitive assessment scores CDR-SB, ADAS-Cog and MMSE. Both the outcome $\mathbf{Y}$ and the mediators $\mathbf{M}$ are scaled without centering ensuring that the values the $\alpha$ and $\beta$ are the standardized regression coefficients. Tables 3, 4 and 5 summarized all the mediator selection results with their estimated mediation effect for changes in CDR-SB, ADAS-Cog and MMSE scores, respectively. For our proposed method with random forest estimation in path-b (GKMS (RF)), we evaluate the mediator effect using g-formula with numerical Monte Carlo integration as described by Imai et al. (2010), For the other methods, we employed structural equation modeling (SEM) through the lavaan R package (Rosseel, 2012). More details on the estimation of mediation effects are provided in Appendix D.3.